OpenStudy (anonymous):

Line Integral/Conservative Vector Fields:Find the value of the line integral F*dr(Hint: If F is conservative, the integration may be easier on an alternative path. ) F(x,y)=<2xy,x^2> (a) r1(t)= from 0 from 0

OpenStudy (anonymous):

Here's the problem retyped \[F(x,y)=<2xy,x^2>\] \[r _{1}(t)=<t,t^2>\] from \[0<t<1\] \[r _{2}(t)=<t,t^3>\] from 0<t<1

OpenStudy (anonymous):

the "<" sign should be less than or equal to

OpenStudy (anonymous):

I know that \[r \prime(t)=<1,2t>\]

OpenStudy (anonymous):

^ This is for r1

OpenStudy (anonymous):

I'm unsure of how to get F(t) though. My book says its \[F(t)=<2t^3,t^2>\]

OpenStudy (anonymous):

Line integral Formula \[\int\limits_{}^{}F*dr\] (retyped this formula that was in my question)

OpenStudy (anonymous):

Maybe you should use the hint.

OpenStudy (anonymous):

F is conservative \[\frac {dM} {dy} = \frac {dN} {dx}=2x\]

OpenStudy (anonymous):

F is conservative, so there is a potential: \[F=\nabla\phi\] So you're looking for phi(x,y) such that dphi/dx= 2xy and dphi/dy=x^2

OpenStudy (anonymous):

Can you compute phi?

OpenStudy (anonymous):

I think so: \[\Delta \phi = <2x,2x> = v\]

OpenStudy (anonymous):

Do you understand the notation? phi is a scalar function. \[\nabla \phi\]is the gradient of phi. so \[\nabla \phi = \left(\begin{matrix}\frac{d\phi}{dx} \\ \frac{d \phi}{dy}\end{matrix}\right)\]

OpenStudy (anonymous):

\[\Delta\]is different from \[\nabla\]

OpenStudy (anonymous):

^ I couldn't figure out how to make that sign on here, I didn't see an option for it.

OpenStudy (anonymous):

\nabla

OpenStudy (anonymous):

^ okay thanks.

OpenStudy (anonymous):

With a good guess you can find phi pretty easily here. Do you see that \[\phi = x^2y\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and taking j's partial derivative of course would be 0 so \[f(x,y)= x^2y\]

OpenStudy (anonymous):

Ah wait, I meant the integral of j would be 2xy

OpenStudy (anonymous):

What is j, and what is f?

OpenStudy (anonymous):

we'll, we know that if the curl of that is 0, then it'll be conservative... so we compute for dQ/dx-dP/dy, where F=<P,Q>=<2xy,x^2> and if this is 0 then it'll be conservative.. so dQ/dx=d (x^2)/dx=2x dP/dy=d (2xy)/dy=2x so dQ/dx-dP/dy=2x-2x=0.. now we have a conservative vector field. now we find the potential function f for F, such that ∇f=F=<2xy,x^2> so we know that \[\int\limits_{}^{}Pdx=\int\limits 2xy dx=f\] \[\int\limits Q dy =\int\limits x^2 dy=f\] I'll first use the 1st equation so: \[f=\int\limits 2xy dx=x^2y+h(y)\] I've put h(y), since in this case, since we're integrating with respect to dx, all other expressions with only y's other constants will be considered constants of integration. Now we need to find h(y).. so since df/dy=Q=x^2 , and df/dy=d(x^2y+h(y))/dy=x^2+h'(y), so x^2+h'(y)=x^2 h'(y)=0 integrating h'(y), we'll have h'(y)=0+c=c, so f becomes f=x^2y+c... now why did we want f? recall that \[\int\limits_{t=a}^{t=b}∇f.rds=f(r(b))-f(r(a))\] where r is a vector recall also that if F is a conservative vector field, then there is a function f such that ∇f=F and in this case, our f=x^2y+c. and so we'll have a.) \[\int\limits_{C}^{}F.dr=\int\limits_{t=a}^{t=b}∇f.dr=\int\limits_{0}^{1}∇(x^2y+c).dr=(t^2t^2)|_{0}^{1}=1 \] b.) \[\int\limits\limits_{C}^{}F.dr=\int\limits\limits_{t=a}^{t=b}∇f.dr=\int\limits\limits_{0}^{1}∇(x^2y+c).dr=(t^2t^3)|_{0}^{1}=1\]

OpenStudy (anonymous):

Sorry my book has a different notation from wikipedia, allow me to make some changes: \[v= \nabla \phi \] is \[F = \nabla f\] in my book and from what you were doing I was assuming that you were trying tofind the potential function for F (or V according to wikipedia) so \[\nabla f(x,y)\] would be \[\nabla v(x,y)\]

OpenStudy (anonymous):

^ that was to Thomas

OpenStudy (anonymous):

ok, so \[f=x^2y\]

OpenStudy (anonymous):

Now you can use the formula \[\int_{r_1} F dr\]=f(endpoint)-f(starting point)

OpenStudy (anonymous):

Is it clear so far?

OpenStudy (anonymous):

anonymous, where did the two t^2's come from?

OpenStudy (anonymous):

the path is (t,t^2), so when you substitute that into f you'll get t^2t^2.

OpenStudy (anonymous):

exactly @thomas9

OpenStudy (anonymous):

oh okay I see now

OpenStudy (anonymous):

I get everything that you are saying but I'm still confused as to how they are getting this for F: \[r \prime (t) = <t,2t> \] \[F(t)=<2t^3,t^2>\] \[\int\limits_{C}^{}F*dr=\int\limits_{0}^{1}=4t^3dt = 1\]

OpenStudy (anonymous):

^ Not the integration part just F(t)

OpenStudy (anonymous):

Oh wait, I think I figured that out

OpenStudy (anonymous):

I think they just substituted t and t^2 int F(t) \[F(t,t^2)=2t^2+t^2\]

OpenStudy (anonymous):

I don't understand your last line, but yes they substituted t for x and t^2 for y in F. Of course that's a completely different approach than what we did with the potential.

OpenStudy (anonymous):

That was from my book, I think the reason why its different was because they took a different approach. even though there's seems shorter, I understood the other method. I think I can solve the second part now. Thanks.

OpenStudy (anonymous):

@anonymoustwo44 I've put h(y), since in this case, since we're integrating with respect to dx, all other expressions with only y's other constants will be considered constants of integration. Now we need to find h(y).. so since \[\frac{df}{dy}=Q=x^2\] and \[\frac{df}{dy}=\frac{d(x^2y+h(y))}{dy=x^2+h^ \prime (y)}\] so \[x^2+h^ \prime (y) = x^2\] \[h^\prime(y)=0\] integrating h'(y), we'll have h'(y)=0+c=c, so f becomes f=x^2y+c... now why did we want f? recall that ^ I had to rewrite what you put down because it was confusing to read at first.

OpenStudy (anonymous):

@anonymoustwo44 Its been a while since I've looked at the problem could you explain to me again how you are getting this part: dfdy=d(x2y+h(y))dy=x2+h′(y)

OpenStudy (anonymous):

ah ok. so I got a f(x,y) from integrating 2xy with respect to dx right? so we got f(x,y)=x^2y+h(y) so how do I find h(y) now. Notice that if I differentiate this with respect to y, it will be equal to df/dy=Q=x^2 so d(x^2y+h(y))/dy=x^2+h'(y) now I also know that df/dy=Q=x^2 so x^2+h'(y)=x^2 h'(y)=0 h(y)=c so f(x,y) becomes: x^2y+c<answer

OpenStudy (anonymous):

Ok thanks for the clarification.